OpenStudy (anonymous):

Answer "then" explanition after or together Will Fan and Metal :D Two equations are given below: a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)? (−2, −2) (−3, −1) (−9, −7) (−5, −3)

Vocaloid (vocaloid):

rewrite the first equation a = 4 + 3b then we have two equations a = 4 + 3b a = b - 2 4 + 3b = b - 2 solve for b

OpenStudy (anonymous):

Sorry I dobt understand

OpenStudy (anonymous):

dont~

Vocaloid (vocaloid):

4 + 3b = b - 2 what is b?

OpenStudy (anonymous):

B= -3

OpenStudy (anonymous):

So is the answer B?

Vocaloid (vocaloid):

no, B is not the answer we know that b = -3 a = b - 2 = ?

OpenStudy (anonymous):

I dont know

Vocaloid (vocaloid):

b = -3 a = b - 2 = ?

Vocaloid (vocaloid):

replace "b" with "-3" and find a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A= -5

OpenStudy (anonymous):

The answer is D

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

thank you I fanned and gave you the metal